Assume that one student in a class of 27 students is randomly selected to win a prize. would it be unusal for that student to win? explain.
Answer by stanbon(75887)

Assume that one student in a class of 27 students is randomly selected to win a prize. would it be unusual for that student to win? explain.
----
Binomial distribution with n = 27 and p = 1/27 =0.037
------
mean = n*p = 1
std = sqrt(npq) = sqrt(1*26/27) = 0.9813
------------
Find the z-value of 1/27
z(0.037) = (0.037-1)/0.9813 = -0.9813
So 1/27 is 0.9813 standard deviations below the mean.
---
Conclusion:: Since 1/27 is not at least 2 standard deviatioins
below the mean, the student's selection is not that unusual.
-----
Cheers,
Stan H.
--------------